Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined overhead rate was based on a cost formula that estimates $1,040,000 of total manufacturing overhead for an estimated activity level of 80,000 machine-hours.

During the year, a large quantity of furniture on the market resulted in cutting back production and a buildup of furniture in the company’s warehouse. The company’s cost records revealed the following actual cost and operating data for the year:

Machine-hours 71,000
Manufacturing overhead cost $ 990,000
Inventories at year-end:
Raw materials $ 11,000
Work in process (includes overhead applied of $73,840) $ 156,000
Finished goods (includes overhead applied of $147,680) $ 312,000
Cost of goods sold (includes overhead applied of $701,480) $ 1,482,000

Required:

1. Compute the underapplied or overapplied overhead.

2. Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry.

3. Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry.

4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold?
Answer #1

1) Overhead rate = 1040000/80000 = 13 per labor hour

Applied overhead = 71000*13 = 923000

Actual overhead = 990000

Under applied overhead = 923000-990000 = 67000

2) Journal entry

Date account and explanation Debit Credit
Cost of goods sold 67000
Manufacturing overhead 67000

3) Journal entry

Date account and explanation Debit Credit
Work in process (67000*8%) 5360
Finished goods 10720
Cost of goods sold 50920
Manufacturing overhead 67000

4) net operating income increase by (5360+10720) = 16080